OpenStudy (anonymous):

(dx/x^p) lower limit of integration 0 and upper limit of integration is 1. Find value of p which the integral converges and evaluate the integral for those value at p.

OpenStudy (lgbasallote):

\[\int_0^1 \frac{dx}{x^p}\] is this the question @mandonut

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

so do you have an idea how to do this?

OpenStudy (anonymous):

i got the integral which is 1/(1-p)(x^(p-1))

OpenStudy (lgbasallote):

hmm

OpenStudy (lgbasallote):

actually you need to turn this into \[\large \int x^{-p} dx\] then use power rule. is that what you did?

OpenStudy (anonymous):

yes

OpenStudy (lgbasallote):

then your power rule should be \[\frac{1}{1 - p} \cdot x^{-p + 1}\]

OpenStudy (lgbasallote):

or for organization purposes \[\large \frac{x^{1-p}}{1-p}\]

OpenStudy (anonymous):

which is the same as my answer if u factor -1 for the exponent of x

OpenStudy (lgbasallote):

oh wait. x(p-1) is in the denominator?

OpenStudy (lgbasallote):

then it's right

OpenStudy (lgbasallote):

x^(p-1) *

OpenStudy (lgbasallote):

so did you try plugging in the limits?

OpenStudy (anonymous):

when 0 its undefined thats what im stuck on

OpenStudy (lgbasallote):

actually when it's 1 then it's undefined

OpenStudy (anonymous):

u plug 1 into x, not p

OpenStudy (lgbasallote):

oops lol yeah

OpenStudy (lgbasallote):

got confused

OpenStudy (lgbasallote):

wait...how did it become undefined?

OpenStudy (anonymous):

1/0^(p-1) and p cannot be equal to 1

OpenStudy (lgbasallote):

that is why you should write it as \[\frac{x^{1-p}}{1-p}\]

OpenStudy (lgbasallote):

see if x = 0 then it's just 0...as lng as p is not 1

OpenStudy (anonymous):

good point

OpenStudy (anonymous):

thanks man

OpenStudy (lgbasallote):

you're welcome

OpenStudy (valpey):

It is undefined for that integral, but not because\[\frac{x^{1-1}}{1-1}\text{ is undefined, it is because}\] \[\int_0^1\frac{dx}{x} = ln(x)\vert_0^1 = \ln(1)-\ln(0)\text{ and }\ln(0)\text{ is undefined}\]

OpenStudy (valpey):

\[lim_{\epsilon\rightarrow 0}\int_\epsilon^1\frac{dx}{x^p}\] For p=1:\[lim_{\epsilon\rightarrow 0}\int_\epsilon^1\frac{dx}{x}=lim_{\epsilon\rightarrow 0}(\ln{x}\vert_{\epsilon}^1)=lim_{\epsilon\rightarrow 0}(\ln{1}-\ln{\epsilon})=0-lim_{\epsilon\rightarrow 0}(\ln{\epsilon})=\]\[0-(-\infty)=\infty\]For p<>1\[lim_{\epsilon\rightarrow 0}\int_\epsilon^1\frac{dx}{x^p}=lim_{\epsilon\rightarrow 0}\frac{x^{1-p}}{1-p}\vert_{\epsilon}^1\]\[=lim_{\epsilon\rightarrow 0}\frac{1}{1-p}-\frac{\epsilon^{1-p}}{1-p}=\frac{1}{1-p}-lim_{\epsilon\rightarrow 0}\frac{\epsilon^{1-p}}{1-p}=\frac{1}{1-p}(1-lim_{\epsilon\rightarrow 0}\epsilon^{1-p})\] This limit only exists when 1-p is positive. Therefore p<1 and the integral = \[\frac{1}{1-p}\]